OpenStudy (moongazer):

If 3^y = z, then 3^(y+2) = Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

3^(y+2) = 3^y * 3^2 = z * 9

OpenStudy (moongazer):

please explain

OpenStudy (anonymous):

(a^x ) (a^y) = a^(x+y)

OpenStudy (anonymous):

u can write that in reverse 2

OpenStudy (moongazer):

sorry i can't understand it that much please explain more

OpenStudy (moongazer):

w8 i think i understand it

OpenStudy (anonymous):

\[so \ 3^{y+2} \ can\ be \ written \ as (3^{y})(3^{2}))\]

OpenStudy (moongazer):

give another sample question and i'll answer it

OpenStudy (anonymous):

5^x = z , then whts 5^(x+3)

OpenStudy (moongazer):

875

OpenStudy (anonymous):

howcome?

OpenStudy (moongazer):

because 5^(x+3) is also equal to (5^x)(5^3) then i replaced the 5^x to z and it will become z(5^3) =125z oooooops. my earlier answer is wrong thought z is 7

OpenStudy (moongazer):

you still there?

OpenStudy (anonymous):

hi im here, are u there?